# Patterns in the Transition From Founder-Leadership to Community Governance of Open Source

### v3

@@ 1.1. Transitions to community governance in open-source software (OSS)
Fortunately, OSS projects seem to agree with this assessment. It is now commonplace that the Internet’s most prominent OSS projects are undergoing transformations in their governance from the founder-owner default to community governance. After six years of operating under an ideologically aligned founder, but without formal community representation, the Debian operating system in 1997 initiated a multi-year governance transition process that concluded in annual democratic elections (which, themselves, continued to evolve in form and scope for several years) (O’mahony and Ferraro, 2007). The Django web framework moved from a single-leader governance model to a core team model in 2014 (Holovaty, 2014) and from a core team model to a more inclusive and community-driven “technical advisory board” in 2020 (Bennett et al., 2020). The developers of the Python programming language, after decades under a single-leader governance model, took inspiration from the Django transition (Warsaw et al., 2018) to implement an elected board structure. In social media, the growth of the early community-led news website Slashdot motivated several redesigns of the site’s moderation system, from manual moderation by the founder to ever-expanding methods of delegating the task to users, ultimately giving reputable users considerable opportunities for effective voice over Slashdot’s content (Ganley and Lampe, 2009). Other prominent examples of OSS projects that have transitioned from founder-owners to community governance include the FreeBSD operating system and the Apache server. At the same time, notable projects such as the Linux operating system and Vim text editor have persisted under founder-owner leadership.
@@ 1.2. Governance documents as a window into institutional processes
Fortunately, as a reflection of its resourcefulness, the OSS community has developed a loose standard for defining OSS project governance on GitHub, one that prescribes a clear, uniform location and format for governing documents. Increasingly, influential OSS projects on GitHub—such as node.js, Docker, and Jupyter—are posting public GOVERNANCE.md files that list out their policies for contributions, donations, planning, and other essential OSS governance functions. GOVERNANCE.md files are an elegant use of the “git” version control software to assign responsibilities and provide clear community entry points into OSS governance (Izquierdo and Cabot, 2015a). In this it follows other emergent conventions for community management, like CONTRIBUTING.md and CODE_OF_CONDUCT.md files. Also, because GOVERNANCE.md is a plain text file co-hosted with project code, it can leverage git’s tools for version control and forking. This means tracking of historical changes by default, and excellent tools for projects to learn from each other’s governance designs, and select their own institutional designs in an informed manner. GOVERNANCE.md files are an entry point for OSS projects to transition from the implicit feudalism of founder-owners to true, and truly accountable community governance.
@@ 1.4. The study
To address the gap in our understanding of governance transitions from unitary to community governance, we compile a dataset of 637 open-source repositories to compare the initial and latest snapshots of the GOVERNANCE.md constitutions, enabling a longitudinal examination of how governance is explicitly articulated and evolves across projects (Yin et al., 2022)(Izquierdo and Cabot, 2015b).We parse text from these artifacts into three core components—Roles, Actions, and Deontics—capturing who is responsible, what activities are authorized or required, and the normative force behind those directives, reflecting established practices for articulating governance rules and governance-oriented roles in OSS ecosystems (Izquierdo and Cabot, 2015b)(Izquierdo and Cabot, 2021). The methodological contribution lies in integrating NLP-driven information extraction with statistical approaches to scale OSS governance analysis across hundreds of projects: we leverage textual embeddings and grouping techniques to surface governance patterns from many repositories, a direction supported by prior work that combines linguistic analyses of OSS communications with longitudinal network perspectives to study governance dynamics (Yin et al., 2022)(Izquierdo and Cabot, 2015b).This approach operationalizes governance as a textual, machine-readable artifact and enables cross-project comparability, consistent with the literature’s emphasis on explicit governance documentation and the use of computational methods to analyze governance-related text in relevant contexts (Izquierdo and Cabot, 2015b)(Yin et al., 2022).
@@ 2. Related Work
Importantly, governance encompasses more than technical coordination; it is a social process that structures participation, trust, and accountability in diverse, geographically dispersed collaboration networks, with group awareness, norms, and transparent decision-making playing central roles in sustaining coordination (Gutwin et al., 2004). Empirical analyses across OSS projects emphasize that governance mechanisms—such as conflict management, contribution guidelines, and social practices that facilitate knowledge sharing—are essential to long-term resilience, not merely auxiliary to code development (Feller et al., 2005)(Vasilescu et al., 2014)(Crowston et al., 2008). Consequently, OSS governance emerges as a socio-technical concern that not only warrants scholarly attention but also raises the need for approaches that can directly capture how governance is articulated and evolves within projects (Germonprez et al., 2018)(Gutwin et al., 2004).
@@ 4. Methods
Figure 1. Processing pipeline from raw governance files to structured institutional statements and analysis. These steps support measurement of change in count and concentration from initial to latest versions of version-controlled GOVERNANCE.md project constitutions. This diagram shows how governance text is normalized, parsed into roles, actions, deontics, and objects, and clustered into institutional constructs. :
@@
Normalization and alignment. We preprocess governance documents by removing badges and images, converting tables to lists, normalizing headings, and stripping markup. Text we segmented into short paragraph blocks and sentences using a splitter tuned for Markdown lists. To reduce pronoun ambiguity, we applied coreference resolution while maintaining a reversible mapping to original offsets (Lee et al., 2018) (Jurafsky and Martin, [n. d.]). Where headings were detectable, we recorded section counts across snapshots to capture the degree of governance structuring.
@@ 5. Results
Figure 2. Projects diversify the range of governance roles over time. Plots show the share of role mentions in initial versus latest governance snapshots. Early constitutions are dominated by broad categories such as “all_project” and “all_community,” while later constitutions redistribute attention across more specialized roles (e.g., subcommittees, technical committees, and steering groups). This broadening reflects institutional development toward greater specialization and shared governance. The differences in the distributions of these types are small but significant.:
Figure 3. Projects expand the catalog of governance actions over time. Plot shows the share of action mentions in initial versus latest governance snapshots. While high-level categories such as “choice” and “authority” remain prominent, later constitutions show a broader and more balanced distribution across action types, reflecting increased institutional complexity and scope. As above, the differences in the distributions of these types are small but significant.:
Figure 4. Deontic composition in OSS governance remains broadly stable over time, although entropy of binomial enabling/restricting distribution increases. Panel A show the share of modal expressions (can/may,” must/will,” should”) in initial versus latest governance snapshots. The relative balance between permissive, obligatory, and advisory language changes little, suggesting that while projects diversify roles and actions, the prescriptive force of their rules remains largely constant. Panel B highlights the distribution of enabling versus restricting deontic statements. Enabling language (“can”, “may”) accounts for over 97% of references in both periods, while restricting terms (“cannot”, “must not”) remain a small minority, and seem to decline further over time. Together, these patterns underscore the relative stability of governance constitutions, where permissions dominate over prohibitions.:
@@
Figure 5. The number of types of roles and actions invoked in constitutions increases significantly. Violin plots compare the number of categories of (a) Roles, (b) Actions, and (c) Deontics between initial and latest snapshots for each repository. Roles and actions both show significant increases.:

### v4

@@ 1.1. Transitions to community governance in open-source software (OSS)
Fortunately, OSS projects seem to agree with this assessment. It is now commonplace that the Internet’s most prominent OSS projects are undergoing transformations in their governance from the founder-owner default to community governance. After six years of operating under an ideologically aligned founder, but without formal community representation, the Debian operating system in 1997 initiated a multi-year governance transition process that concluded in annual democratic elections (which, themselves, continued to evolve in form and scope for several years) (O’mahony and Ferraro, 2007). The Django web framework moved from a single-leader governance model to a core team model in 2014 (Holovaty, 2014) and from a core team model to a more inclusive and community-driven “technical advisory board” in 2020 (Bennett and others, 2020). The developers of the Python programming language, after decades under a single-leader governance model, took inspiration from the Django transition (Warsaw and others, 2018) to implement an elected board structure. In social media, the growth of the early community-led news website Slashdot motivated several redesigns of the site’s moderation system, from manual moderation by the founder to ever-expanding methods of delegating the task to users, ultimately giving reputable users considerable opportunities for effective voice over Slashdot’s content (Ganley and Lampe, 2009). Other prominent examples of OSS projects that have transitioned from founder-owners to community governance include the FreeBSD operating system and the Apache server. At the same time, notable projects such as the Linux operating system and Vim text editor have persisted under founder-owner leadership.
@@ 1.2. Governance documents as a window into institutional processes
Fortunately, as a reflection of its resourcefulness, the OSS community has developed a loose standard for defining OSS project governance on GitHub, one that prescribes a clear, uniform location and format for governing documents. Increasingly, influential OSS projects on GitHub—such as node.js, Docker, and Jupyter—are posting public GOVERNANCE.md files that list out their policies for contributions, donations, planning, and other essential OSS governance functions. GOVERNANCE.md files are an elegant use of the “git” version control software to assign responsibilities and provide clear community entry points into OSS governance (Izquierdo and Cabot, 2015b). In this it follows other emergent conventions for community management, like CONTRIBUTING.md and CODE_OF_CONDUCT.md files. Also, because GOVERNANCE.md is a plain text file co-hosted with project code, it can leverage git’s tools for version control and forking. This means tracking of historical changes by default, and excellent tools for projects to learn from each other’s governance designs, and select their own institutional designs in an informed manner. GOVERNANCE.md files are an entry point for OSS projects to transition from the implicit feudalism of founder-owners to true, and truly accountable community governance.
@@ 1.4. The study
To address the gap in our understanding of governance transitions from unitary to community governance, we compile a dataset of 637 open-source repositories to compare the initial and latest snapshots of the GOVERNANCE.md constitutions, enabling a longitudinal examination of how governance is explicitly articulated and evolves across projects (Yin et al., 2022)(Izquierdo and Cabot, 2015a).We parse text from these artifacts into three core components—Roles, Actions, and Deontics—capturing who is responsible, what activities are authorized or required, and the normative force behind those directives, reflecting established practices for articulating governance rules and governance-oriented roles in OSS ecosystems (Izquierdo and Cabot, 2015a)(Izquierdo and Cabot, 2021). The methodological contribution lies in integrating NLP-driven information extraction with statistical approaches to scale OSS governance analysis across hundreds of projects: we leverage textual embeddings and grouping techniques to surface governance patterns from many repositories, a direction supported by prior work that combines linguistic analyses of OSS communications with longitudinal network perspectives to study governance dynamics (Yin et al., 2022)(Izquierdo and Cabot, 2015a).This approach operationalizes governance as a textual, machine-readable artifact and enables cross-project comparability, consistent with the literature’s emphasis on explicit governance documentation and the use of computational methods to analyze governance-related text in relevant contexts (Izquierdo and Cabot, 2015a)(Yin et al., 2022).
@@ 2. Related Work
Importantly, governance encompasses more than technical coordination; it is a social process that structures participation, trust, and accountability in diverse, geographically dispersed collaboration networks, with group awareness, norms, and transparent decision-making playing central roles in sustaining coordination (Gutwin et al., 2004). Empirical analyses across OSS projects emphasize that governance mechanisms—such as conflict management, contribution guidelines, and social practices that facilitate knowledge sharing—are essential to long-term resilience, not merely auxiliary to code development (Feller et al., 2005)(Vasilescu et al., 2014)(Crowston et al., 2008). Consequently, OSS governance emerges as a socio-technical concern that not only warrants scholarly attention but also raises the need for approaches that can directly capture how governance is articulated and evolves within projects (Noori et al., 2025)(Germonprez et al., 2018)(Gutwin et al., 2004).
@@ 4. Methods
Figure 1. Processing pipeline from raw governance files to structured institutional statements and analysis. These steps support measurement of change in count and concentration from initial to latest versions of version-controlled GOVERNANCE.md project constitutions. This diagram shows how governance text is normalized, parsed into roles, actions, and deontics, and clustered into institutional constructs.
A left-to-right block diagram showing ingestion of GOVERNANCE.md files, parsing into roles/actions/deontics, clustering, and metrics (count and entropy) feeding result plots.
@@
Normalization and alignment. We preprocess governance documents by removing badges and images, converting tables to lists, normalizing headings, and stripping markup. Text we segmented into short paragraph blocks and sentences using a splitter tuned for Markdown lists. To reduce pronoun ambiguity, we applied coreference resolution while maintaining a reversible mapping to original offsets (Lee et al., 2018) (Jurafsky and Martin, ). Where headings were detectable, we recorded section counts across snapshots to capture the degree of governance structuring.
@@ 5. Results
Figure 2. Projects diversify the range of governance roles over time. Plots show the share of role mentions in initial versus latest governance snapshots. Early constitutions are dominated by broad categories such as “all_project” and “all_community,” while later constitutions redistribute attention across more specialized roles (e.g., subcommittees, technical committees, and steering groups). This broadening reflects institutional development toward greater specialization and shared governance. The differences in the distributions of these types are small but significant.
Figure 3. Projects expand the catalog of governance actions over time. Plot shows the share of action mentions in initial versus latest governance snapshots. While high-level categories such as “choice” and “authority” remain prominent, later constitutions show a broader and more balanced distribution across action types, reflecting increased institutional complexity and scope. As above, the differences in the distributions of these types are small but significant.
Figure 4. Deontic composition in OSS governance remains broadly stable over time, although entropy of binomial enabling/restricting distribution increases. Panel A show the share of modal expressions (can/may,” must/will,” should”) in initial versus latest governance snapshots. The relative balance between permissive, obligatory, and advisory language changes little, suggesting that while projects diversify roles and actions, the prescriptive force of their rules remains largely constant. Panel B highlights the distribution of enabling versus restricting deontic statements. Enabling language (“can”, “may”) accounts for over 97% of references in both periods, while restricting terms (“cannot”, “must not”) remain a small minority, and seem to decline further over time. Together, these patterns underscore the relative stability of governance constitutions, where permissions dominate over prohibitions.
@@
Figure 5. The number of types of roles and actions invoked in constitutions increases significantly. Violin plots compare the number of categories of (a) Roles, (b) Actions, and (c) Deontics between initial and latest snapshots for each repository. Roles and actions both show significant increases.
Three side-by-side violin plots. The first shows role distributions for initial and latest snapshots, the second shows action distributions, and the third shows deontic distributions; each violin reflects per-repository values across the corpus.